# CaRtGS: Computational Alignment for Real-Time Gaussian Splatting SLAM

### v2

@@ Abstract
Simultaneous Localization and Mapping (SLAM) is pivotal in robotics, with photorealistic scene reconstruction emerging as a key challenge. To address this, we introduce Computational Alignment for Real-Time Gaussian Splatting SLAM (CaRtGS), a novel method enhancing the efficiency and quality of photorealistic scene reconstruction in real-time environments. Leveraging 3D Gaussian Splatting (3DGS), CaRtGS achieves superior rendering quality and processing speed, which is crucial for scene photorealistic reconstruction. Our approach tackles computational misalignment in Gaussian Splatting SLAM (GS-SLAM) through an adaptive strategy that optimizes training, addresses long-tail optimization, and refines densification. Experiments on Replica and TUM-RGBD datasets demonstrate CaRtGS’s effectiveness in achieving high-fidelity rendering with fewer Gaussian primitives. This work propels SLAM towards real-time, photorealistic dense rendering, significantly advancing photorealistic scene representation. For the benefit of the research community, we release the code on our project website: https://dapengfeng.github.io/cartgs.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) is a cornerstone of robotics and has been a subject of extensive research over the past few decades [1, 2, 3]. The rapid evolution of applications such as autonomous driving, virtual and augmented reality, and embodied intelligence has introduced new challenges that extend beyond the traditional scope of real-time tracking and mapping. Among these challenges is the need for photorealistic scene reconstruction, which necessitates precise spatial understanding coupled with high-fidelity visual representation.
In response to these challenges, recent research has explored the use of implicit volumetric scene representations, notably Neural Radiance Fields (NeRF) [4]. While promising, integrating NeRF into SLAM systems has encountered several obstacles, including high computational demands, lengthy training times, limited generalizability, an over-reliance on visual cues, and a susceptibility to catastrophic forgetting [5].
@@
However, existing Gaussian Splatting SLAM (GS-SLAM) methods [7, 8, 9, 10, 11, 12, 13, 14, 15] struggle to achieve superior rendering performance under real-time constraints when dealing with a limited number of Gaussian primitives. These issues stem from the misalignment between the computational demands of the algorithm and the available processing resources, which can lead to insufficient training and optimization processes. Addressing these challenges is crucial for enhancing the performance and applicability of GS-SLAM in real-time environments.
@@
  • We provide an in-depth analysis of the computational misalignment phenomenon present in GS-SLAM.
  • We introduce an adaptive computational alignment strategy that effectively tackles insufficient training, long-tail optimization, and weak-constrained densification, achieving high-fidelity rendering with fewer Gaussian primitives under real-time constraints.
  • We conduct comprehensive experiments and ablation studies to demonstrate the effectiveness of our proposed method on two popular datasets with two distinct camera types.
@@ II Related Works
GS-SLAM leverages the benefits of 3DGS [6] to achieve enhanced performance in terms of rendering speed and photorealism. LABEL:Fig:Schema illustrates the foundational structure of GS-SLAM, which is based on a 3D Gaussians Map and encompasses four key functional modules: localization, geometric mapping, photorealistic rendering, and loop closure. In this section, we conduct a concise review of both 3D Gaussian Splatting and Gaussian Splatting SLAM.
@@
Motivated by these advancements, our work addresses the challenge of insufficient training in photorealistic rendering within real-time SLAM by utilizing splat-wise backpropagation [18]. In parallel, recent methodologies have concentrated on sparse-view reconstruction and have sought to compact the scene representation. This is achieved by training a neural network to serve as a data-driven prior, which is capable of directly outputting Gaussians in a single forward pass [22, 23, 24, 25]. In contrast, our research zeroes in on real-time dense-view and per-scene visual SLAM. This targeted focus demands an incremental photorealistic rendering output that is tailored to the unique characteristics of each scene.
(a) Schema
(b) Performance on TUM-RGBD
Figure 1: The overview of Gaussian Splatting SLAM. (a) The schema of Gaussian Splatting SLAM. (b) We provide a comprehensive comparison of most of the available open-source GS-SLAM methods using the TUM-RGBD dataset, captured with an RGBD camera.
@@ II-B Gaussian Splatting SLAM
These methods achieve state-of-the-art PSNR with a large number of Gaussian primitives, as presented in LABEL:Fig:Performance-on-TUM-RGBD. This phenomenon is attributed to the computational misalignment that occurs during photorealistic rendering in GS-SLAM. In this paper, we delve into this phenomenon and propose an innovative computational alignment technique to enhance PSNR while reducing the number of Gaussian points required, all within the constraints of real-time SLAM operations.
@@ III Methods
In this section, we delve into the photorealistic rendering aspect of GS-SLAM. Initially, we scrutinize the computational misalignment phenomenon inherent to GS-SLAM, as detailed in III-A. This misalignment can significantly impair computational efficiency and hinder the swift convergence of photorealistic rendering, adversely affecting the performance of real-time GS-SLAM. To overcome these obstacles, we propose a novel adaptive computational alignment strategy, presented in III-B. This strategy aims to accelerate the 3DGS process, optimize computational resource allocation, and efficiently control model complexity, thereby enhancing the overall effectiveness and practicality of 3DGS in real-time SLAM applications.
@@ III-A Computational Misalignment
The computational misalignment encountered in photorealistic rendering within the context of SLAM can be attributed to three primary aspects: insufficient training, long-tail optimization, and weak-constrained densification, which reduces rendering quality and increases map size. These factors significantly hinder the real-time applications of GS-SLAM, limiting its applicability in resource-constrained devices.
Figure 2: The Effect of Adaptive Optimization. The dashed line indicates the results obtained without the application of our adaptive optimization technique, whereas the solid line represents the results when our optimization is employed. Notably, our adaptive optimization significantly improves the processing of keyframes that exhibit a low PSNR by enhancing the iterative optimization process.
III-A1 Insufficient Training
In contrast to typical 3DGS [6], which is not constrained by real-time considerations, online rendering within the realm of SLAM necessitates the concurrent execution of localization, mapping, and rendering at a speed that is synchronized with the frequency of incoming sensor data. To achieve this, the majority of current real-time GS-SLAM methods [13, 15, 14] rely on keyframes for both mapping and rendering. However, these methods typically achieve only a few thousand iterations in rendering optimization in total, which significantly lags behind the tens of thousands of iterations achieved by 3DGS [6]. Due to insufficient training, the optimization process has not fully converged, adversely affecting the quality of online rendering.
Recent observations by several researchers indicate that pixel-wise backpropagation in 3DGS presents a significant computational challenge [19, 18]. This process becomes a bottleneck due to the contention among multiple GPU threads for access to shared Gaussian primitives, which necessitates serialized atomic operations, thereby limiting parallelization efficiency. Unfortunately, this drawback is integrated into the previous implementations of GS-SLAM [13, 15, 14]. In this paper, we utilize a fast splat-wise backpropagation [18] to reduce thread contention. This approach not only achieves a 3×3$\times$3 × increase in the number of iterations compared to the baseline [15], but also maintains the same runtime. This advancement significantly mitigates the problem of insufficient training, substantially improving the rendering quality of real-time GS-SLAM.
Figure 3: The overview of CaRtGS. We adopt ORB-SLAM3 as a front-end tracker, severing for localization and geometry mapping. In the photorealistic rendering back-end, we apply the proposed adaptive computational alignment strategy to enhance the 3DGS optimization process, including fast splat backward, adaptive optimization, and opacity regularization.
@@ III-A2 Long-Tail Optimization
To mitigate the issue of catastrophic forgetting, a common approach in GS-SLAM is to randomly select a keyframe from the keyframe pool for periodic retraining[15, 13, 14]. However, this method can result in suboptimal long-tail optimization, as depicted in LABEL:Fig:adaptive-opt. Specifically, the retraining frequency of the earliest keyframes tends to exceed that of the most recently added ones. This disparity arises because the keyframe pool is continuously expanded as the camera moves through the environment, which can result in an uneven distribution of retraining efforts and and a declining trend in the PSNR for newly incoming keyframes.
@@ III-A3 Weak-constrained Densification
Densification is a critical component of photorealistic rendering in the context of GS-SLAM, encompassing both geometry densification and adaptive densification[8, 12, 9, 11, 13, 7, 15, 14, 10]. Geometric densification involves the conversion of a colored point cloud into initialized Gaussian primitives for each newly identified keyframe, providing a foundational geometric structure for the environment. Adaptive densification, on the other hand, refines the Gaussian primitives using operations such as splitting and cloning, which are guided by gradients and the size of the primitives themselves [6]. These densifications are solely constrained by a simplistic pruning strategy that eliminates Gaussian primitives with low opacity. However, emerging research [23, 24, 25] suggests that this approach is insufficient for managing the model’s size within an optimal range. In this paper, we introduce an opacity regularization loss to encourage the Gaussian primitives to learn a low opacity, thereby not only facilitating the pruning process to eliminate less significant primitives but also preserving high-fidelity rendering.
III-B Adaptive Computational Alignment
An overview of our system is delineated in LABEL:Fig:Overview-of-CaRtGS. Building upon the cutting-edge advancements in real-time GS-SLAM [15], we adopt ORB-SLAM3 [3] as the front-end tracker, which provides not only efficient 6-DoF camera pose estimations of the input image $\mathcal{V}_{i}$, but also colored point cloud $\mathcal{P}$. In the localization component, the front-end tracker interatively refine the camera orientation $\mathbf{R}\in SO(3)$ and position $\mathbf{t}\in\mathbb{R}^{3}$ by minimizing the reprojection error between 2D ORB keypoint $\mathbf{p}_{i}$ and matched 3D point $\mathbf{P}_{i}$ with Levenberg–Marquardt (LM) algorithm:
[…]
where $\Sigma_{g}$ is the covariance matrix associated with $\mathbf{p}_{i}$ and $\mathbf{P}_{i}$, $\pi\left(\cdot\right)$ represents the projection function, and $\rho$ is the Huber cost function. In the geometry mapping component, the front-end tracker performs a local bundle adjustment on a set of co-visible 3D points $\mathcal{P}_{L}$ and keyframes $\mathcal{K}_{L}$ to generate a colored point cloud $\mathcal{P}_{i}$.
@@ III-B Adaptive Computational Alignment
Given a colored point cloud $\mathcal{P}_{i}$, we transform it into a set of initialized Gaussian primitives $\mathcal{G}_{i}$. Each primitive is characterized by its posistion $\mathbf{p}\in\mathbb{R}^{3}$, orientation represented as quaternion $\mathbf{q}\in\mathbb{R}^{4}$ , scaling factor $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\sigma\in\mathbb{R}^{1}$, and spherical harmonic coefficients $\mathbf{SH}\in\mathbb{R}^{48}$. Subsequently, we merge $\mathcal{G}_{i}$ into the 3D Gaussians Map $\mathcal{G}$ through the geometry densification operation. Utilizing 3DGS[6], we can get the high-fidelity rendering $\hat{\mathcal{I}}$ for a given 6-DoF camera pose:
^$\hat{\mathcal{I}}=\sum_{i}c_{i}\alpha_{i}\prod_{j=1}^{i-1}\left(1-\alpha_{i}\right),$ (2)
where $c_{i}$ denotes the color derived from $\mathbf{SH}$, $\alpha_{i}$ is determined by evaluating a projected 2D Gaussian multipied with the learned opacity $\sigma_{i}$.
@@
III-B1 Fast Splat-wise Backpropagation
@@
In the conventional 3DGS training pipeline, the backpropagation phase is computationally demanding as it entails the propagation of gradient information from pixels to Gaussian primitives. This process necessitates the calculation of gradients for each splat-pixel pair $(i,j)$, followed by an aggregation step. In each iteration, GPU thread $i+1$ applies the standard $\alpha$-blending logic to transition from the received state $\mathcal{X}_{i,j}$ to $\mathcal{X}_{i+1,j}$, integrating this updated information into the gradient computation. This process can be mathematically represented as:

### v4

@@ Abstract
Simultaneous Localization and Mapping (SLAM) is pivotal in robotics, with photorealistic scene reconstruction emerging as a key challenge. To address this, we introduce Computational Alignment for Real-Time Gaussian Splatting SLAM (CaRtGS), a novel method enhancing the efficiency and quality of photorealistic scene reconstruction in real-time environments. Leveraging 3D Gaussian Splatting (3DGS), CaRtGS achieves superior rendering quality and processing speed, which is crucial for scene photorealistic reconstruction. Our approach tackles computational misalignment in Gaussian Splatting SLAM (GS-SLAM) through an adaptive strategy that enhances optimization iterations, addresses long-tail optimization, and refines densification. Experiments on Replica, TUM-RGBD, and VECtor datasets demonstrate CaRtGS’s effectiveness in achieving high-fidelity rendering with fewer Gaussian primitives. This work propels SLAM towards real-time, photorealistic dense rendering, significantly advancing photorealistic scene representation. For the benefit of the research community, we release the code and accompanying videos on our project website: https://dapengfeng.github.io/cartgs.
Index Terms:
Mapping, Gaussian Splatting SLAM, SLAM.
@@ I Introduction
Simultaneous Localization and Mapping (SLAM) is a cornerstone of robotics and has been a subject of extensive research over the past few decades [1, 2, 3, 4, 5]. The rapid evolution of applications such as autonomous driving, virtual and augmented reality, and embodied intelligence has introduced new challenges that extend beyond the traditional scope of real-time tracking and mapping. Among these challenges is the need for photorealistic scene reconstruction, which necessitates precise spatial understanding coupled with high-fidelity visual representation.
In response to these challenges, recent research has explored the use of implicit volumetric scene representations, notably Neural Radiance Fields (NeRF) [6]. While promising, integrating NeRF into SLAM systems has encountered several obstacles, including high computational demands, lengthy optimization times, limited generalizability, an over-reliance on visual cues, and a susceptibility to catastrophic forgetting [7].
@@
However, existing Gaussian Splatting SLAM (GS-SLAM) methods [9, 10, 11, 12, 13, 14, 15, 16, 17] struggle to achieve superior rendering performance under real-time constraints when dealing with a limited number of Gaussian primitives. These issues stem from the misalignment between the computational demands of the algorithm and the available processing resources, which can lead to insufficient optimization and optimization processes. Addressing these challenges is crucial for enhancing the performance and applicability of GS-SLAM in real-time environments.
@@
  • We provide an analysis of the computational misalignment phenomenon present in GS-SLAM.
  • We introduce an adaptive computational alignment strategy that effectively tackles insufficient optimization, long-tail optimization, and weak-constrained densification, achieving high-fidelity rendering with fewer Gaussian primitives under real-time constraints.
  • We conduct comprehensive experiments and ablation studies to demonstrate the effectiveness of our proposed method on three popular datasets with three distinct camera types.
@@ II Related Works
GS-SLAM leverages the benefits of 3DGS [8] to achieve enhanced performance in terms of rendering speed and photorealism. In this section, we conduct a concise review of both 3D Gaussian Splatting and Gaussian Splatting SLAM.
@@
Motivated by these advancements, our work addresses the challenge of insufficient optimization in photorealistic rendering within real-time SLAM by utilizing splat-wise backpropagation [20]. In parallel, recent methodologies have concentrated on sparse-view reconstruction and have sought to compact the scene representation. This is achieved by training a neural network to serve as a data-driven prior, which is capable of directly outputting Gaussians in a single forward pass [24, 25, 26, 27]. In contrast, our research zeroes in on real-time dense-view and per-scene visual SLAM. This targeted focus demands an incremental photorealistic rendering output that is tailored to the unique characteristics of each scene.
Figure 1: Performance on TUM-RGBD. We provide a comparison of most of the available open-source GS-SLAM methods.
@@ II-B Gaussian Splatting SLAM
These methods achieve state-of-the-art PSNR with a large number of Gaussian primitives, as presented in LABEL:Fig:Overview-of-GS-SLAM, which will limit the application of real-time GS-SLAM in large-scale scenarios due to increased computational demands. In this paper, we delve into the limitations of existing GS-SLAM and propose an innovative computational alignment technique to enhance PSNR while reducing the number of Gaussian primitives required, all within the constraints of real-time SLAM operations.
@@ III Methods
In this section, we delve into the photorealistic rendering aspect of GS-SLAM. Initially, we scrutinize the computational misalignment phenomenon inherent to GS-SLAM. This misalignment can significantly impair computational efficiency and hinder the swift convergence of photorealistic rendering, adversely affecting the performance of real-time GS-SLAM. To overcome these obstacles, we propose a novel adaptive computational alignment strategy. This strategy aims to accelerate the 3DGS process, optimize computational resource allocation, and efficiently control model complexity, thereby enhancing the overall effectiveness and practicality of 3DGS in real-time SLAM applications.
@@ III-A Computational Misalignment
The computational misalignment encountered in photorealistic rendering within the context of SLAM can be attributed to three primary aspects: insufficient optimization, long-tail optimization, and weak-constrained densification, which reduces rendering quality and increases map size. These factors significantly hinder the real-time applications of GS-SLAM, limiting its applicability in resource-constrained devices.
Figure 2: The Effect of Adaptive Optimization on Replica. Dashed lines depict performance without adaptive optimization, while solid lines show results with it. Blue represents keyframe iterations, and red indicates PSNR. The horizontal line marks average PSNR and iterations. Our method significantly improves low-PSNR keyframe processing through enhanced iterative optimization, as evident from the trend comparison between dashed and solid lines.
III-A1 Insufficient Optimization
In contrast to typical 3DGS [8], which is not constrained by real-time considerations, online rendering within the realm of SLAM necessitates the concurrent execution of localization, mapping, and rendering at a speed that is synchronized with the frequency of incoming sensor data. To achieve this, the majority of current real-time GS-SLAM methods [15, 17, 16] rely on keyframes for both mapping and rendering. However, these methods typically achieve only a few thousand iterations in rendering optimization in total, which significantly lags behind the tens of thousands of iterations achieved by 3DGS [8]. Due to insufficient optimization, the optimization process has not fully converged, adversely affecting the quality of online rendering.
Recent observations by several researchers indicate that pixel-wise backpropagation in 3DGS presents a significant computational challenge [21, 20]. This process becomes a bottleneck due to the contention among multiple GPU threads for access to shared Gaussian primitives, which necessitates serialized atomic operations, thereby limiting parallelization efficiency. Unfortunately, this drawback is integrated into the previous implementations of GS-SLAM [15, 17, 16]. In this paper, we utilize a fast splat-wise backpropagation [20] to reduce thread contention. This approach not only achieves a 3×3$\times$3 × increase in the number of iterations compared to the baseline [17], but also maintains the same runtime. This advancement significantly mitigates the problem of insufficient optimization, substantially improving the rendering quality of real-time GS-SLAM.
Figure 3: The overview of CaRtGS. We adopt a real-time cutting-edge SLAM system as a front-end tracker, severing for localization and geometry mapping. In the photorealistic rendering back-end, we apply the proposed adaptive computational alignment strategy to enhance the 3DGS optimization process, including fast splat backward, adaptive optimization, and opacity regularization.
@@ III-A2 Long-Tail Optimization
To mitigate the issue of catastrophic forgetting, a common approach in GS-SLAM is to randomly select a keyframe from the keyframe pool for periodic reoptimization[17, 15, 16]. However, this method can result in suboptimal long-tail optimization, which overfits the oldest keyframe and underfits the newest one, as depicted in LABEL:Fig:adaptive-opt. Specifically, the reoptimization frequency of the earliest keyframes tends to exceed that of the most recently added ones. This disparity arises because the keyframe pool is continuously expanded as the camera moves through the environment, which can result in an uneven distribution of reoptimization efforts and and a declining trend in the PSNR for newly incoming keyframes.
@@ III-A3 Weak-constrained Densification
Densification is a critical component of photorealistic rendering in the context of GS-SLAM, encompassing both geometry densification and adaptive densification[10, 14, 11, 13, 15, 9, 17, 16, 12]. Geometric densification involves the conversion of a color point cloud into initialized Gaussian primitives for each newly identified keyframe, providing a foundational geometric structure for the environment. Adaptive densification, on the other hand, refines the Gaussian primitives using operations such as splitting and cloning, which are guided by gradients and the size of the primitives themselves [8]. These densifications are solely constrained by a simplistic pruning strategy that eliminates Gaussian primitives with low opacity. However, emerging research [25, 26, 27] suggests that this approach is insufficient for managing the model’s size within an optimal range. In this paper, we introduce an opacity regularization loss to encourage the Gaussian primitives to learn a low opacity, thereby not only facilitating the pruning process to eliminate less significant primitives but also preserving high-fidelity rendering.
III-B System Overview
As delineated in LABEL:Fig:Overview-of-CaRtGS, we take the modular designs, which are easy to integrate into existing real-time decoupled GS-SLAM, e.g., GS-ICP-SLAM[16] and Photo-SLAM [17].
Given a sequence of observations $\{\mathcal{V}_{1},...,\mathcal{V}_{N}\}$, we employ a state-of-the-art front-end tracker [2, 3], which estimates the 6-DoF pose for each frame and identifies keyframes $\{v_{1},...,v_{k}\}$ based on criteria related to translation and rotation. Once a keyframe $v_{i}$ is identified, the frontend tracker transforms the corresponding observation $\mathcal{V}_{i}$ into the global coordinate system and integrates it into the global Point Cloud $\mathcal{P}$.
@@ III-B System Overview
In the photorealistic rendering phase, we utilize 3DGS [8] as the backend render. Firstly, we convert $\mathcal{P}$ into a set of Gaussian primitives $\mathcal{G}$. Each primitive is characterized by its posistion $\mathbf{p}\in\mathbb{R}^{3}$, orientation represented as quaternion $\mathbf{q}\in\mathbb{R}^{4}$ , scaling factor $\mathbf{s}\in\mathbb{R}^{3}$, opacity $\sigma\in\mathbb{R}^{1}$, and spherical harmonic coefficients $\mathbf{SH}\in\mathbb{R}^{48}$. By employing $\alpha$–blending rendering [8], we achieve the high-fidelity rendering $\hat{\mathcal{I}}$ for a selected keyframe $v_{i}$:
^$\hat{\mathcal{I}}=\sum_{k\in\mathcal{G}}c_{k}\alpha_{k}\prod_{j=1}^{k-1}\left(1-\alpha_{k}\right),$ (1)
where $c_{k}$ denotes the color derived from $\mathbf{SH}$, $\alpha_{k}$ is determined by evaluating a projected 2D Gaussian multipied with the learned opacity $\sigma_{k}$. To refine the Gaussian primitives $\mathcal{G}$, we take both $\mathcal{L}_{1}$ and Structural Similarity Index (SSIM) Loss $\mathcal{L}_{ssim}$ to supervise the optimization process. These losses are crucial for enhancing the quality of our photorealistic renderings. Additionally, we incorporate opacity regularization into our comprehensive loss function to control the model size, which is detailed in Sec. LABEL:Subsec:Opacity-Regularization.
III-C Adaptive Computational Alignment
@@
III-C1 Fast Splat-wise Backpropagation
@@
In the conventional 3DGS optimization pipeline, the backpropagation phase is computationally demanding as it entails the propagation of gradient information from pixels to Gaussian primitives. This process necessitates the calculation of gradients for each splat-pixel pair $(i,j)$, followed by an aggregation step. In our notation, $i$ denotes the index of the $i$-th splat, and $j$ denotes the index of the $j$-th pixel. To parallelize the execution, we assign thread $i$ to process the $i$-th splat, and thread $j$ to process the $j$-th pixel. In the forward pass, GPU thread $i+1$ applies the standard $\alpha$-blending logic to transition from the received state $\mathcal{X}_{i,j}$ to $\mathcal{X}_{i+1,j}$, integrating this updated information into the gradient computation. In the backward pass, the gradients associated with the $i$-th splat, denoted as $\nabla\mathcal{X}_{i}$, are accumulated across the pixels that are influenced by this splat. This process can be mathematically represented as: